OpenStudy (anonymous):

What is the vertex of the graph of y = |–x – 3|+ 1?

OpenStudy (anonymous):

y = |–x – 3|+ 1 y = |–x – 3|+ 1 y=|-(x+3)| + 1 y=|-(x-(-3))| + 1 compare with y=|x-h| + k vertex is given by (h,k) so vertex (h,k)=(-3,1)